OpenStudy (anonymous):

Which of the equations represents the function? A) y = 4x2 B) y = -4x2 C) y = 1 4 x2 D) y = 1 2 x2

OpenStudy (anonymous):

uhh could you please rewrite c and d

OpenStudy (cwrw238):

what are the last 2 y = (1/4)x^2 y = (1/2) x^2 ?

OpenStudy (anonymous):

A. y=4x^2 B. y=-4x^2 C. y=1/4x^2 D. y=1/2x^2 sorry my computer does that sometimes

OpenStudy (cwrw238):

from graph when y =5, x = +- 4.5 or close 1/4 (4.5^2) = approx 5 so its C

OpenStudy (anonymous):

yes cwrw is right

OpenStudy (anonymous):

ok thanks!

OpenStudy (cwrw238):

yw